1. Aggregate demand curve of an economy is given by AD = 51 - 0.2P, the long-run aggregate supply, LRAS, is 30 and the short-run aggregate supply is given by SRAS = 0.3 P (all output measures are in US$ billions and the price level is given as an index number). What could be the unemployment rate if the natural rate of unemployment is 4%?

2. Aggregate demand curve of an economy is given by AD = 51 - 0.2P, the long-run aggregate supply, LRAS, is 30 and the short-run aggregate supply is given by SRAS = 0.3 P (all output measures are in US$ billions and the price level is given as an index number). Assume that the present status of the economy is the result of a demand shock. What should be the original price level when the economy was in equilibrium

3. Aggregate demand curve of an economy is given by AD = 51 - 0.2P, the long-run aggregate supply, LRAS, is 30 and the short-run aggregate supply is given by SRAS = 0.3 P (all output measures are in US$ billions and the price level is given as an index number). Assume that the present status of the economy is the result of a demand shock. What should be the price level when the labor market adjusts and the economy will be in an equilibrium?
Answer #1

1. Aggregate demand curve of an economy is given by AD = 51 - 0.2P, the long-run aggregate supply, LRAS, is 30 and the short-run aggregate supply is given by SRAS = 0.3 P (all output measures are in US$ billions and the price level is given as an index number). What could be the unemployment rate if the natural rate of unemployment is 4%

in short run, SRAS = AD, gives P = 102

at p = 102, Yt = 0.3*102 = 30.6

therefore, Yt- Yn = 30.6 - 30 = 0.6

un = 4, so by okun's law, ut - un = (-)a.(Yt-Yn)

not knwoing the value of a, we have ut = (-)a.(Yt-Yn) + un = -0.6a + 4

ut = -0.6a + 4

Aggregate demand curve of an economy is given by AD = 51 - 0.2P, the long-run aggregate supply, LRAS, is 30 and the short-run aggregate supply is given by SRAS = 0.3 P (all output measures are in US$ billions and the price level is given as an index number). Assume that the present status of the economy is the result of a demand shock. What should be the original price level when the economy was in equilibrium

in equilibrium, AD = LRAS

51 - 0.2P = 30

51-30 = 0.2P

P = 21/0.2 = 105